# Generating Exceptional Behavior Tests with Reasoning Augmented Large Language Models

### v2

@@ Abstract
Many popular programming languages, including C#, Java, and Python, support exceptions. Exceptions are thrown during program execution if an unwanted event happens, e.g., a method is invoked with an illegal argument value. Software developers write exceptional behavior tests (EBTs) to check that their code detects unwanted events and throws appropriate exceptions. Prior research studies have shown the importance of EBTs, but those studies also highlighted that developers put most of their efforts on “happy paths”, e.g., paths without unwanted events. To help developers fill the gap, we present the first framework, dubbed exLóng, that automatically generates EBTs. exLóng is a large language model instruction-tuned from CodeLlama and embeds reasoning about traces that lead to throw statements, conditional expressions that guard throw statements, and non-exceptional behavior tests that execute similar traces. We compare exLóng with the state-of-the-art models for test generation (CAT-LM) and one of the strongest foundation models (GPT3.5), as well as with analysis-based tools for test generation (Randoop and EvoSuite). Our results show that exLóng outperforms existing models and tools. Furthermore, we contributed several pull requests to open-source projects and 23 EBTs generated by exLóng were already accepted.
@@ 5 ByteArrayInputStream ins = new ByteArrayInputStream(bytes);
((b)) Exceptional behavior test for the highlighted statement above.
@@
Sadly, tool support for automatically generating EBTs is limited. Most existing analysis-based test generation tools (e.g., Randoop [32, 38] and EvoSuite [12]) and learning-based test generation tools (e.g., CAT-LM [36] and TeCo [29]) have no special settings for targeting EBTs and are primarily evaluated on non-EBTs. Random test generation tools can be guided by reinforcement learning to target exceptional behaviors [1], but the generation works only on the entire code base, and not for a specific throw statement that a developer might select. Additionally, tests produced by analysis-based tools lack readability [7, 33, 6].
@@
We present the first framework, dubbed exLóng, an instruction-tuned large language model (LLM) that automatically generates EBTs. LLMs are shown to be effective in code generation, including test generation [52, 36, 29, 23, 59]. Such strong prior provides a good foundation yet is not enough. EBTs contribute to only a very small percentage in existing code bases, i.e., they are not well-represented in LLM training data. The reasoning trajectory that trigger an EBT is uncommon and not captured by the Causal Language Modeling pretraining objective [5, 35], and thus LLMs trained on typical (non-EBT) code does not perform well on the task of generating EBTs.
@@
Using CodeLlama [39] as its foundation, exLóng is instruction fine-tuned [56, 58, 41] with a novel task instruction and fine-tuning data, designed specifically to embed the reasoning about a context that includes: (a) traces that lead to target throw statements, (b) guard expressions (i.e., conditional expressions that guard those throw statements), and (c) non-EBTs that execute similar traces. This context is used as the input to generate an EBT that triggers the target throw statement. The EBT that we already showed in Figure 1(b) was generated by exLóng.
@@
We assess the power of exLóng using two use cases. In the first use case, which we call developer-oriented use case, a developer selects a method under test and a target throw statement, as well as a destination test file. exLóng takes these inputs and automatically generates an EBT that executes the target throw statement.
@@
In this use case, we compare exLóng with the state-of-the-art models for test generation (CAT-LM [36]) and one of the strongest foundation models (GPT3.5 [5, 31]). We use a number of standard metrics (BLEU [34], CodeBLEU [37], edit similarity [60, 45] and exact match), as well as metrics specific to code, including percentage of compilable tests, executable tests, and executable tests that cover the target throw statement. Our results show that exLóng generates 68.6% and 10.9% more executable EBTs than CAT-LM and GPT3.5, respectively.
@@
In the second use case, which we call machine-oriented use case, a developer uses exLóng to automatically generate EBTs for the entire code base with the goal to cover all existing throw statements (with one EBT per statement). exLóng takes the entire code base as input, finds throw statements that are in public methods already covered by at least one non-EBT and generates one EBT for each of the throw statements. This use case is similar to the traditional test generation setup targeted by analysis-based generation tools.
@@
In this use case, we compare exLóng with popular analysis-based test generation tools: Randoop [32, 38] and EvoSuite [12]. Although tools complement each other (i.e., each tool can generate EBTs for some target throw statements that other tools cannot), our findings show that exLóng outperforms Randoop and EvoSuite.
@@
Finally, we selected a subset of EBTs generated by exLóng and created pull requests for several open-source projects. By the time of this writing, 23 tests generated by exLóng have already been accepted by developers of those projects.
@@
  • Model. We designed and implemented exLóng, an instruction-tuned LLM built on CodeLlama, which reasons about traces to methods that contain throw statements, guard expressions, and non-EBTs that cover similar traces.
  • Use cases. We recognized two use cases for exLóng: developer- and machine-oriented use cases.
  • Evaluation. We assess the power of exLóng in both use cases. In developer-oriented use case, we compare exLóng with existing models for code and test generation. In machine-oriented use case, we compare exLóng with analysis-based testing tools: Randoop and EvoSuite. We find that exLóng outperforms existing state-of-the-art models and tools.
  • Dataset. We developed a novel dataset for the presented task. We plan to make this dataset publicly available upon acceptance of this paper.
II Task
Figure 2: Overview of exLóng. Two use cases for exLóng: (1) developer-oriented use case and (2) machine-oriented use case. In the developer-oriented use case, a developer specify the method under test, a target throw statement and a destination test file and ask exLóng to generate an EBT that cover the target throw statement. In the machine-oriented use case, a developer gives an entire repository to exLóng.
At a high level, exLóng is designed to help software developers write EBTs that cover the throw statements within the given repository. We propose two use cases for exLóng: developer-oriented use case (Section II-A) and machine-oriented use case (Section II-B). Note that in this work we do not consider generating EBTs that cover throw statements in the dependency libraries of the given repository, e.g. $\mathtt{ArithmeticException}$ thrown from the $\mathtt{java.lang.Math}$, as we assume that those throw statements are already covered by EBTs available in dependency libraries; developers can configure exLóng to include such throw statements if necessary.
@@ II-A Developer-oriented use case
In the developer-oriented use case (box 1 in Figure 2), a developer will specify the method under test (MUT), a target throw statement, and a destination test file, then ask exLóng to generate an EBT that invokes MUT and triggers the target throw statement. When the developer specifies a target throw statement that is not within the MUT, exLóng will first use a static program analysis technique to find all possible throw statements reachable through a sequence of method calls starting from the MUT, and then prompt the developer to select the target statement from the list.
@@ II-B Machine-oriented use case
In the machine-oriented use case (box 2 in Figure 2), a developer gives an entire repository to exLóng. exLóng will first find public methods and throw statements in those methods. Furthermore, exLóng will heuristically select a destination test file. Then, exLóng will create (public method, throw statement, destination test file) triples. Finally, it will generate one EBT to test each triple, without requiring any developer intervention. exLóng uses static+dynamic program analyses to obtain necessary context such as traces leading to throw statements and guard expressions.
III exLóng
Figure 2 shows the workflow of exLóng. Given an MUT ($\mathtt{m_{ut}}$), a target throw statement ($\mathtt{s}$), and a destination test file ($\mathtt{d}$), exLóng reasons about the following context, collected using static and dynamic program analyses ( 3):
@@
  • relevant non-EBTs ($\hat{\mathtt{T}}_{\mathtt{neb}}$): a subset of non-EBTs whose snippets and coding styles may help with generating EBTs for the given inputs.
The above constitutes the prompt $\mathtt{p}=(\mathtt{m_{ut}},\mathtt{s},\mathtt{d},\mathtt{r},\mathtt{g},\hat{\mathtt{T}}_{\mathtt{neb}})$ that encompasses the task inputs and the task context ( 4). During training, exLóng is instruction-tuned over a foundation LLM to produce EBT ($\mathtt{t}_{\mathtt{eb}}$) conditioned on the input $\mathtt{p}$. During inference, exLóng generates the EBT ( 6) given the prompt.
While the format of the LLM’s input ($\mathtt{p}$) and output ($\mathtt{t}_{\mathtt{eb}}$) is static, the steps for preparing the task inputs and reasoning about the task context differ for training and inference, and differ slightly between the developer- and machine-oriented use cases. We describe the methodology for training exLóng in Section III-A and inference in Section III-B.
@@ III-A Training
1:input: $\mathtt{T}_{\mathtt{eb}}$, $\mathtt{T}_{\mathtt{neb}}$ existing EBTs and non-EBTs
@@
3:output: $\mathcal{C}$ the training corpus of $(\mathtt{m_{ut}},\mathtt{s},\mathtt{d},\mathtt{r},\mathtt{g},\hat{\mathtt{T}}_{\mathtt{neb}},\mathtt{t}_{\mathtt{eb}})$}
@@
To instruction-tune exLóng’s foundation LLM, we collect a corpus ($\mathcal{C}$) of instruction data $\mathtt{c}=(\mathtt{p},\mathtt{t}_{\mathtt{eb}})=(\mathtt{m_{ut}},\mathtt{s},\mathtt{d},\mathtt{r},\mathtt{g},\hat{\mathtt{T}}_{\mathtt{neb}},\mathtt{t}_{\mathtt{eb}})$ from a set of repositories with developer-written EBTs and non-EBTs. Algorithm 1 shows the process of collecting the training corpus for exLóng.
@@ III-A2 Executing EBT and collecting stack trace
Each EBT will be expanded to one instruction data ($\mathtt{c}$) in the training corpus ($\mathcal{C}$). Naturally, the file that contains the EBT is the destination test file (line 7 in Algorithm 1). To avoid data leakage problems, we remove all test methods in the test file ($\mathtt{d}$) and only keep the test class structure and utility methods.
@@ III-A3 Computing the guard expression
Stack trace provides the sequence of method invocations that lead to the target throw statement, but knowing only the names of the methods is insufficient for generating EBTs. To aid the reasoning about the setup of the system under test, which lead to exceptional behaviors, we propose guard expression: a logical formula representing the constraints on the symbolic variables that must be true to follow the particular code trace. Specifically, we use conjunctions of expressions extracted from the invoked methods in the stack trace to form the guard expression. For instance, the guard expression for the throw statement highlighted in Figure 1(a) is present in the Figure 2 (box 3).
The first step in the computation is to collect the list of guard-related AST nodes along the stack trace starting from the target throw statement to the MUT, as described in Algorithm 2. We traverse each method in the stack trace in reversed order (line 5 in Algorithm 2). Inside each method, we start from the AST node specified by the line number in the stack trace. We always include this node to the list of collected nodes (line 8 in Algorithm 2) because the variable names in that statement may be used by the next step for replacing method call arguments. Starting from the statement, we traverse the AST by maintaining the pointer $\mathtt{current}$ that constantly moves from the child AST node to its parent node. AST nodes entailing condition expressions within the ‘for’ loop (lines 10 to 11 in Algorithm 2), ‘if’ statement (lines 12 to 16 in Algorithm 2), ‘while’ loop and ‘switch’ statement will be added to the set $\mathtt{N}$. We also collect the assignment statements, method call expressions, and method declarations. They are omitted from Algorithm 2 due to space constraints.
The second step of computing guard expressions is to process the collected nodes by propagating the symbolic variables through the stack trace, as described in Algorithm 3. The nodes are visited in the order being collected, and conditional expressions are directly added to the guard expression (line 9 in Algorithm 3). Assignment statements, method declaration, and method call expressions will trigger a Merge operation (line 21 in Algorithm 3). The goal of the Merge operation is to review the current guard expression and replace the symbolic variables (e.g., variables appearing in the target throw statement) with their corresponding values (e.g., constant values or MUT’s arguments). Therefore, the guard expression reflects the MUT’s arguments and public fields that are usable by EBT, rather than local variables that EBT may not have access to. For assignment statement (line 11 in Algorithm 3), we replace the left-hand side variable in $\mathcal{E}$ with the right hand side expression. For method declaration and method call expression (lines 15 to 18 in Algorithm 3; these two nodes always appear in pairs in the collected nodes), we replace the method declaration’s argument names in $\mathcal{E}$ with the actual arguments in the method call expression.

### v3

# exLong: Generating Exceptional Behavior Tests with Large Language Models

@@ Abstract
Many popular programming languages, including C#, Java, and Python, support exceptions. Exceptions are thrown during program execution if an unwanted event happens, e.g., a method is invoked with an illegal argument value. Software developers write exceptional behavior tests (EBTs) to check that their code detects unwanted events and throws appropriate exceptions. Prior research studies have shown the importance of EBTs, but those studies also highlighted that developers put most of their efforts on “happy paths”, e.g., paths without unwanted events. To help developers fill the gap, we present the first framework, dubbed exLong, that automatically generates EBTs. exLong is a large language model instruction fine-tuned from CodeLlama and embeds reasoning about traces that lead to throw statements, conditional expressions that guard throw statements, and non-exceptional behavior tests that execute similar traces. We compare exLong with the state-of-the-art models for test generation (CAT-LM) and one of the strongest foundation models (GPT-4o), as well as with analysis-based tools for test generation (Randoop and EvoSuite). Our results show that exLong outperforms existing models and tools. Furthermore, we contributed several pull requests to open-source projects and 23 EBTs generated by exLong were already accepted.
Index Terms:
test generation, large language models, program analysis, exceptional behavior tests
@@ 5 ByteArrayInputStream ins = new ByteArrayInputStream(bytes);
((b)) Exceptional behavior test written using JUnit 4 that covers the highlighted statement above.
@@
Sadly, tool support for automatically generating EBTs is limited. Most existing analysis-based test generation tools (e.g., Randoop [36, 42] and EvoSuite [11]) and learning-based test generation tools (e.g., CAT-LM [40] and TeCo [32]) have no special settings for targeting EBTs and are primarily evaluated on non-EBTs. Random test generation tools can be guided by reinforcement learning to target exceptional behaviors [1], but the generation works only on the entire codebase, and not for a specific throw statement that a develop might select. Additionally, tests produced by analysis-based tools lack readability [6, 37, 5].
@@
We present the first framework, dubbed exLong, an instruction fine-tuned large language model (LLM) that automatically generates EBTs. LLMs are shown to be effective in code generation, including test generation [56, 40, 32, 22, 63]. Such strong prior provides a good foundation yet is not enough. EBTs contribute to only a very small percentage in existing codebases, i.e., they are not well-represented in LLM training data. The special conditions that trigger an EBT during execution are not included in the training phase of standard code LLMs, thus they do not perform well on the task of generating EBTs.
@@
Using CodeLlama [43] as its base, exLong is fine-tuned [60, 62, 45] with a novel task instruction and fine-tuning data, designed specifically to embed the reasoning about a context that includes: (a) traces that lead to target throw statements, (b) guard expressions (i.e., conditional expressions that guard those throw statements), and (c) non-EBTs that execute similar traces. This context is used as the input to generate an EBT that triggers the target throw statement. The EBT that we already showed in Figure 1(b) was generated by exLong.
@@
We assess the power of exLong using two use cases. In the first use case, which we call developer-oriented use case, a developer selects a method under test and a target throw statement, as well as a destination test file. exLong takes these inputs and automatically generates an EBT that executes the target throw statement.
@@
In this use case, we compare exLong with the state-of-the-art models for test generation (CAT-LM [40]) and strongest foundation models (GPT3.5 [34] and GPT-4o [35]). We use a number of standard metrics (BLEU [39], CodeBLEU [41], edit similarity [64, 49] and exact match), as well as metrics specific to code, including percentage of compilable tests, executable tests, and executable tests that cover the target throw statement. Our results show that exLong generates 83.8% and 9.9% more executable EBTs than CAT-LM and GPT3.5, respectively.
@@
In the second use case, which we call machine-oriented use case, a developer uses exLong to automatically generate EBTs for the entire codebase with the goal to cover all existing throw statements (with one EBT per statement). exLong takes the entire codebase as input, finds throw statements that are in public methods already covered by at least one non-EBT and generates one EBT for each of the throw statements. This use case is similar to the traditional test generation setup targeted by analysis-based generation tools.
@@
In this use case, we compare exLong with popular analysis-based test generation tools: Randoop [36, 42] and EvoSuite [11]. Although tools complement each other (i.e., each tool can generate EBTs for some target throw statements that other tools cannot), our findings show that exLong outperforms Randoop and EvoSuite.
Additionally, we built exLong on GPT-4o (a state-of-the-art language model) without fine-tuning and evaluated it in developer-oriented use case. Our results show that exLong–GPT-4o outperforms GPT-4o by up to 16.6%. This emphasizes that our technique is generalizable to the most advanced proprietary LLMs.
@@
Finally, we selected a subset of EBTs generated by exLong and created pull requests for several open-source projects. By the time of this writing, 23 tests generated by exLong have already been accepted by developers of those projects.
@@
  • Model. We designed and implemented exLong, an instruction fine-tuned LLM built on CodeLlama, which reasons about traces to methods that contain throw statements, guard expressions, and non-EBTs that cover similar traces.
  • Use cases. We recognized two use cases for exLong: developer- and machine-oriented use cases.
  • Evaluation. We assess the power of exLong in both use cases. In developer-oriented use case, we compare exLong with existing models for code and test generation. In machine-oriented use case, we compare exLong with analysis-based testing tools: Randoop and EvoSuite. We find that exLong outperforms existing state-of-the-art models and tools.
  • Dataset. We developed a novel dataset for the presented task and this dataset is publicly available.
exLong is available on GitHub at https://github.com/EngineeringSoftware/exLong.
II Use Cases
Figure 2: Overview of exLong. Two use cases for exLong: (1) developer-oriented use case and (2) machine-oriented use case. In the developer-oriented use case, a developer specify the method under test, a target throw statement and a destination test file and ask exLong to generate an EBT that cover the target throw statement. In the machine-oriented use case, a developer gives an entire repository to exLong.
At a high level, exLong is designed to help software developers write EBTs that cover the throw statements within the given repository. We propose two use cases for exLong: developer-oriented use case (Section II-A) and machine-oriented use case (Section II-B). Note that in this work we do not consider generating EBTs that cover throw statements in the dependency libraries of the given repository, e.g., $\mathtt{ArithmeticException}$ thrown from the $\mathtt{java.lang.Math}$, as we assume that those throw statements are already covered by EBTs available in dependency libraries; developers can configure exLong to include such throw statements if necessary.
@@ II-A Developer-oriented use case
In the developer-oriented use case (box 1 in Figure 2), a developer will specify the method under test (MUT), a target throw statement, and a destination test file, then ask exLong to generate an EBT that invokes MUT and triggers the target throw statement. When the developer specifies a target throw statement that is not within the MUT, exLong will first use a static program analysis technique to find all possible throw statements reachable through a sequence of method calls starting from the MUT, and then prompt the developer to select the target statement from the list.
@@ II-B Machine-oriented use case
In the machine-oriented use case (box 2 in Figure 2), a developer gives an entire repository to exLong. exLong will first find all public methods and throw statements in those methods. Furthermore, exLong will heuristically select a destination test file. Then, exLong will create (public method, throw statement, destination test file) triples. Finally, it will generate one EBT to test each triple, without requiring any developer intervention. exLong uses static+dynamic program analyses to obtain necessary context such as traces leading to throw statements and guard expressions.
III exLong
Figure 2 shows the workflow of exLong. Given an MUT ($\mathtt{m_{ut}}$), a target throw statement ($\mathtt{s}$), and a destination test file ($\mathtt{d}$), exLong reasons about the following context, collected using static and dynamic program analyses ( 3):
@@
  • relevant non-EBTs ($\hat{\mathtt{T}}_{\mathtt{neb}}$): a subset of non-EBTs that invoke the given MUT or are in the given destination test file; their snippets and coding styles help with generating desired EBTs.
The above constitutes the prompt $\mathtt{p}=(\mathtt{m_{ut}},\mathtt{s},\mathtt{d},\mathtt{r},\mathtt{g},\hat{\mathtt{T}}_{\mathtt{neb}})$ that encompasses the task inputs and the task context ( 4). During training, exLong is instruction fine-tuned over a base LLM to produce EBT ($\mathtt{t}_{\mathtt{eb}}$) conditioned on the input $\mathtt{p}$. During inference, exLong generates the EBT ( 6) given the prompt.
While the format of the LLM’s input ($\mathtt{p}$) and output ($\mathtt{t}_{\mathtt{eb}}$) is static, the steps for preparing the task inputs and reasoning about the task context differ for training and inference, and differ slightly between the developer- and machine-oriented use cases. We describe the methodology for training exLong in Section III-A and inference in Section III-B.
@@ III-A Training
1:inputs: $\mathtt{T}_{\mathtt{eb}}$, $\mathtt{T}_{\mathtt{neb}}$ existing EBTs and non-EBTs
@@
3:outputs: $\mathcal{C}$ the training corpus of $(\mathtt{m_{ut}},\mathtt{s},\mathtt{d},\mathtt{r},\mathtt{g},\hat{\mathtt{T}}_{\mathtt{neb}},\mathtt{t}_{\mathtt{eb}})$}
@@
To perform supervised fine-tuning (SFT) on exLong’s foundation LLM, we collect a corpus ($\mathcal{C}$) of SFT data $\mathtt{c}=(\mathtt{p},\mathtt{t}_{\mathtt{eb}})=(\mathtt{m_{ut}},\mathtt{s},\mathtt{d},\mathtt{r},\mathtt{g},\hat{\mathtt{T}}_{\mathtt{neb}},\mathtt{t}_{\mathtt{eb}})$ from a set of repositories with developer-written EBTs and non-EBTs. Algorithm 1 shows the process of collecting the training corpus for exLong.
@@ III-A2 Executing EBT and collecting stack trace
Each EBT will be expanded to one SFT example ($\mathtt{c}$) in the training corpus ($\mathcal{C}$). Naturally, the file that contains the EBT is the destination test file (line 7 in Algorithm 1). To avoid data leakage problems, we remove all test methods in the test file ($\mathtt{d}$) and only keep the test class structure and utility methods.
@@ III-A3 Computing the guard expression
Stack trace provides the sequence of method invocations that lead to the target throw statement, but knowing only the names of the methods is insufficient for generating EBTs. To aid the reasoning about the setup of the system under test, which lead to exceptional behaviors, we propose guard expression: a logical formula representing the constraints on the symbolic variables that must be true to follow the particular code trace. Specifically, we use conjunctions of expressions extracted from the invoked methods in the stack trace to form the guard expressions. For instance, the guard expression for the throw statement highlighted in Figure 1(a) is present in the Figure 2 (box 3).
The first step in the computation is to collect the list of guard-related AST nodes along the stack trace starting from the target throw statement to the MUT, as described in Algorithm 2. We traverse each method in the stack trace in reversed order (line 5). Inside each method, we start from the AST node specified by the line number in the stack trace. We always include this node into the list of collected nodes (line 8) because the variable names in that statement may be used by the next step for replacing method call arguments. Starting from the statement, we traverse the AST by maintaining the pointer $\mathtt{current}$ that constantly moves from the child AST node to its parent node. AST nodes entailing condition expressions within the ‘for’ loop (lines 10 to 11), ‘if’ statement (lines 12 to 16), ‘while’ loop and ‘switch’ statement will be added to $\mathtt{N}$. We also collect the assignment statements, method call expressions, and method declarations. Some of these are omitted from Algorithm 2 to keep it simple.
The second step of computing guard expressions is to process the collected nodes by propagating the symbolic variables through the stack trace, as described in Algorithm 3. The nodes are visited in the order being collected, and conditional expressions are directly added to the guard expression (line 9). Assignment statements, method declaration, and method call expressions will trigger a Merge operation (line 21). The goal of the Merge operation is to review the current guard expression and replace the symbolic variables (e.g., variables appearing in the target throw statement) with their corresponding values (e.g., constant values or MUT’s arguments). Therefore, the guard expression reflects the MUT’s arguments and public fields that are usable by EBT, rather than local variables that EBT may not have access to. For assignment statement (line 11), we replace the left hand side variable in $\mathcal{E}$ with the right hand side expression. For method declaration and method call expression (lines 15 to 18; these two nodes always appear in pairs in the collected nodes), we replace the method declaration’s argument names in $\mathcal{E}$ with the actual arguments in the method call expression.